OpenStudy (anonymous):

The factors of a polynomial are (x + 3)(x - 2)(x + 7). The polynomial has been graphed. How do the solutions relate to the factors?

OpenStudy (zale101):

from this factor (x + 3)(x - 2)(x + 7) we know that x + 3=0 x=-3 x - 2=0 x=2 x + 7=0 x=-7 so -3, 2, and -7 are all the x-intercepts of the polynomials